### v1

@@ 2.4 Prediction-Enhanced Monte Carlo: Evaluation
  1. 1.
    A specific parameter set $\bm{\theta}:=(\bm{\theta}_{\text{model}},\bm{\theta}_{\text{simulation}},\bm{\theta}_{\text{payoff}})\in\bm{\Theta}$ is given, e.g., calibrated from market data in real-time, and needs to be used for pricing. These parameters can be determined through combination of calibration, customization, client agreement or external observation.
  2. 2.
    Generate $n$ pairs of $(\bm{Y}_{i},\bm{X}_{i})_{i\in[n]}$, from $\mathbb{E}_{\text{risk neutral measure}(\bm{\theta})}$. Here we have suppressed $\bm{Y}$’s dependency on $\bm{\theta}$, as in the evaluation stage only a fixed $\bm{\theta}$ is considered.
  3. 3.
    Based on the marginal distribution of $\bm{X}$, independently generate $N$ additional samples of $(\bm{\tilde{X}}_{j})_{j\in[N]}$ from $\mathbb{E}_{\text{risk neutral measure}(\bm{\theta})}$, directly according to its marginal distribution. Importantly, these samples are independent of the previous data $(\bm{Y}_{i},\bm{X}_{i})_{i\in[n]}$.
  4. 4.
@@
    $PEMC:=\frac{1}{n}\sum_{i=1}^{n}(f(\bm{Y}_{i})-g(\bm{X}_{i}))+\frac{1}{N}\sum_{j=1}^{N}g(\bm{\tilde{X}}_{j})$ (4)
For notational simplicity, we have written ${g}(\bm{X})$ instead of ${g}(\bm{\theta}_{\text{model}},\bm{\theta}_{\text{simulation}},\bm{\theta}_{\text{payoff}},\bm{X})$, and $f(\bm{Y})$ instead of $f_{\bm{\theta}_{\text{payoff}}}(\bm{Y})$. A key aspect of this procedure is the relationship between $N$ and $n$. Typically, we choose $N$ to be one, or several orders of magnitude larger than $n$. The optimal choice of $N$ versus $n$, will be discussed Section 3. The PEMC evaluation step is summarized in Algorithm 2.
1:procedure Evaluation($\bm{\theta},n,N,g$)
2:     Generate i.i.d. $(\bm{Y}_{i},\bm{X}_{i})_{i\in[n]}$ from $\mathbb{E}_{\text{risk neutral measure}(\bm{\theta})}$
3:     Generate i.i.d. $(\bm{\tilde{X}}_{j})_{j\in[N]}$ independently from its marginal in $\mathbb{E}_{\text{risk neutral measure}(\bm{\theta})}$
4:     Compute $PEMC:=\frac{1}{n}\sum_{i=1}^{n}(f(\bm{Y}_{i})-g(\bm{X}_{i}))+\frac{1}{N}\sum_{j=1}^{N}g(\bm{\tilde{X}}_{j})$
@@
2.5 PEMC: An Example on Asian Options
To demonstrate the PEMC framework, we return to our ongoing example on arithmetic Asian call option under the Heston model. It is worth noting that while the geometric Asian option is a natural control variate in this context, its closed-form expression is not available under the Heston model. To use control variates here, one must resort to sophisticated methods, e.g, numerically approximating the characteristic function Fusai and Meucci (2008) or employing new framewrok for conditional Monte Carlo Dingec et al. (2015).
This section serves as a user manual, guiding readers through the process of implementing PEMC for pricing an exotic option. When pricing an arithmetic Asian call option under the Heston model, the option payoff is given by:
@@ 2.5 PEMC: An Example on Asian Options
where $\bm{\theta_{\text{payoff}}}$ specifies the strike level $K$, the sampling frequency $n_{D}$ and the observation dates $\{t_{i}\}_{i\in[n_{D}]}$. The Heston model is described in (1) which requires $\bm{\theta}_{\text{model}}:=(r,\eta,\delta,\rho,\kappa)\in\mathbb{R}^{5}$. For this example, let’s say we use a neural network (NN) as the prediction model. Then, the PEMC framework could proceed as follows:
  1. 1.
    Define the parameter space $\bm{\Theta}$ that encompasses realizations of
    $\bm{\theta}:=(\underbrace{r,\eta,\delta,\rho,\kappa}_{\bm{\theta}_{\text{model}}},\underbrace{S_{0},\nu_{0},\Delta t,T}_{\bm{\theta}_{\text{simulation}}},\underbrace{K,n_{D},\{t_{i}\}_{i\in[n_{D}]}}_{\bm{\theta}_{\text{payoff}}})$
    which are of practical interest. The choice of $\bm{\Theta}$ also depends on the model capacity and the model updating frequency. We shall discuss this in subsequent sections.
  2. 2.
    Uniformly sample $\bm{\theta}$ from $\bm{\Theta}$. This step is straightforward if the parameter space is a Cartesian product of intervals. For each sampled $\bm{\theta}$, generate process $\bm{Y}:=(S_{t},\nu_{t})_{t}$ using a Euler scheme with step size $\Delta t$, according to the Heston model specified by $\bm{\theta}$.
  3. 3.
    During the sampling of $\bm{Y}$, we collect $\bm{X}$ to be the sum of Brownian increment
@@
    simulated during each step of the Euler scheme. This makes the marginal of $\bm{X}$ simply a two-dimensional Gaussian: $\frac{\bm{X}}{\sqrt{T}}\sim\mathcal{N}\Big{(}\mathbf{0},[\begin{smallmatrix}1&\rho\\ \rho&1\end{smallmatrix}]\Big{)}$.
  4. 4.
@@
  6. 6.
    Train a NN with weights $\bm{w}$ to minimize the MSE loss:
    $\min_{{\bm{w}}}\frac{1}{N_{\text{train}}}\sum_{i\in[N_{\text{train}}]}(\text{label}_{i}-NN_{\bm{w}}(\text{feature}_{i}))^{2}$
@@
    $NN_{\bm{\hat{w}}}\approx\mathbb{E}\bigg{[}(\frac{1}{n_{D}}\sum_{i=1}^{n_{D}}S_{t_{i}}-K)^{+}\bigg{|}{r,\eta,\delta,\rho,\kappa},{S_{0},\nu_{0},\Delta t,T},{K,n_{D},\{t_{i}\}_{i\in[n_{D}]}},W_{T}^{S},W_{T}^{\nu}\bigg{]}.$
  7. 7.
    At evaluation, a specific $\bm{\theta}=({r,\eta,\delta,\rho,\kappa},{S_{0},\nu_{0},\Delta t,T},{K,n_{D},\{t_{i}\}_{i\in[n_{D}]}},W_{T}^{S},W_{T}^{\nu})\in\Theta$ is given, we generate $n$ paired samples $(\textit{label}_{i},\textit{feature}_{i})_{i\in[n]}$ as in (4). We also generate $N$ i.i.d. $\sqrt{T}\mathcal{N}\Big{(}\mathbf{0},[\begin{smallmatrix}1&\rho\\ \rho&1\end{smallmatrix}]\Big{)}$ samples and store them as $(\tilde{\textit{feature}})_{j\in[N]}$.
  8. 8.
    Output $PEMC:=\frac{1}{n}\sum_{i=1}^{n}(\textit{label}_{i}-NN_{\bm{\hat{w}}}(\textit{feature}_{i}))+\frac{1}{N}\sum_{j=1}^{N}NN_{\bm{\hat{w}}}(\tilde{\textit{feature}}_{j})$
2.6 Implementation Details in PEMC
@@
2.6.1 Parameter Space $\Theta$
@@
When determining the parameter space $\bm{\Theta}$, a key consideration is how frequently the model will be updated. To ensure the performance of PEMC, it is desirable that the training data encompass all practical scenarios the model is expected to encounter within the designated update period. For example, if one is calibrating a Heston model (1) for certain SPY ETF, on a given trading day, the calibration yields a specific set of parameters: $(S_{0},\nu_{0},r,\eta,\delta,\rho,\kappa)=(520,20\%,4.5\%,0.04,0.3,-0.7,0.2)$. If one updates the prediction model in PEMC once a month, then, based on historical data and market conditions, financial practitioners could make reasonable guesses about the bounds within which these parameters are likely to fluctuate over the coming month. For example, one could postulate the underlying price would mostly be fluctuating within a ±$\textpm 5-6\%$ range, i.e., around 490-550, based on the historical volatility of SPY. This expectation specifies an interval for $S_{0}$ in $\bm{\Theta}$. Similarly, one might set intervals for each parameter in the same manner, e.g., interval for $r\in(4.0\%,5.0\%)$, anticipating that the interest rates change would not exceed half a percentage point. Finally, we note that while these parameters are jointly distributed and could be calibrated together—especially considering that many market conditions are intertwined and correlated—this is often unnecessary. Instead, one can define intervals for each parameter individually and enlarge the uncertainty set, since the coverage of practical scenarios is more important in PEMC.
@@
This process is conceptually similar to the selection of uncertainty sets in robust optimization Ben-Tal et al. (2008) or distributionally robust optimization Esfahani and Kuhn (2018), where the goal is to include realizations of parameters with high likelihood. However, unlike robust optimization approaches, which often favor data-dependent uncertainty sets, our focus here is also guided by the expertise of financial engineers. The training parameter space is designed to reflect realistic and practical scenarios derived from domain knowledge, rather than being strictly driven by statistical guarantees. In applications, we construct reasonable $\bm{\Theta}$ for practical models, such as forward curves in HJM modeling Glasserman (2013b) or local stochastic volatility 2D grids Gatheral (2006), among others.
2.6.2 The Design of $\Large{X}$: Feature Engineering
@@
In our Asian option example, we designed the feature $\bm{X}$ as the sum of two correlated Brownian increments. As discussed in Section 2.3, the choice was motivated by two key properties:
@@
  • Predictability: The sum of Brownian motion increments explains away a large portion of the variance driving the movement of SDE.
@@
  • Ease of Generation: The sum of Brownian increments follows a correlated Gaussian distribution, which can be generated efficiently.
Given these reasons, it is natural to use such sums as $\bm{X}$ for all related SDE driving by Brownian motion. Indeed, this is the approach we adopt in our applications, and it has proven to be both robust and effective. For cases where the path is very long, one can divide the sum into chunks of smaller increments. However, we found that using just one increment typically gives sufficient variance reduction.
@@
However, the choice of Brownian increments, while effective, is certainly not the only one. As with most ML tasks, feature engineering that leverages domain expertise can yield better results. We illustrate with one such example, the pricing of a floating strike lookback call option, under the Heston model(1). The option payoff is given by:
@@
This approximation leads to the expression: $f(\{S_{t}\}_{t})\approx S_{0}e^{(r-\nu/2)T+\sqrt{\nu}[\bm{X}]_{1}}-S_{0}e^{[\bm{X}]_{2}}$, which could serve as a potentially effective approximation of $f(\{S_{t}\}_{t})$. This corresponds directly to Property 1 in the construction of $\bm{X}$ in 2.3. For Property 2, the marginal distribution of $\bm{X}$ permits efficient and parallelizable simulation. In particular, applying Girsanov’s Theorem enables one to derive a closed-form expression for the joint density of $([\bm{X}]_{1},[\bm{X}]_{2})$ Karatzas and Shreve (1991). This result makes it straightforward to sample these components directly.
2.6.3 Eliminating Data Storage Overheads
In typical ML workflows, training datasets—often costly, scarce, and carefully curated—are stored and reused extensively. By contrast, within the PEMC framework, the nature of data generation during the prediction model training in Algorithm 1, actually does not require extensive data storage. This difference stems from two main considerations.
First, the volume of training data required to achieve a well-performing prediction model can be large. Storing all of it would be both expensive and unnecessary. In PEMC, the training data can be produced directly via Monte Carlo simulation, ensuring an effectively unlimited supply. Second, this flexibility allows for a more efficient workflow. Data can be generated “on the fly” and processed in streaming fashion. For example, to train on a large number of $N_{\text{train}}$ samples, one could iteratively produce small batches, train the model on these batches, and discard them after training. In this manner, we have successfully trained models in the applications using data on the order of $10^{7}–10^{8}$.
3 Analysis and Performance
In this section, we provide a detailed and comprehensive analysis of the PEMC estimator, presenting both theoretical results and practical considerations. All properties of PEMC discussed here pertain to the evaluation stage, where $\bm{\theta}$ is fixed. Henceforth, we omit the explicit dependence on $\bm{\theta}$ in the notation (e.g., writing $\bm{Y}$ in place of $\bm{Y}(\bm{\theta})$). We begin by establishing the unbiasedness of our estimator.
3.1 Bias Analysis
Theorem 3.1 (Unbiasedness)

### v2

@@ 3.4. Prediction-Enhanced Monte Carlo: Evaluation
  1. (1)
    A specific parameter $\boldsymbol{\theta}:=(\boldsymbol{\theta}_{\text{model}},\boldsymbol{\theta}_{\text{simulation}},\boldsymbol{\theta}_{\text{payoff}})\in\boldsymbol{\Theta}$ is given, e.g., calibrated from market data in real-time, and needs to be used for pricing.
  2. (2)
    Generate $n$ pairs of $(\boldsymbol{Y}_{i},\boldsymbol{X}_{i})_{i\in[n]}$, from $\mathbb{E}_{\text{risk neutral measure}(\boldsymbol{\theta})}$. Here we have suppressed $\boldsymbol{Y}$’s dependency on $\boldsymbol{\theta}$, as in the evaluation stage only one fixed $\boldsymbol{\theta}$ is considered.
  3. (3)
    Based on the marginal distribution of $\boldsymbol{X}$, independently generate $N$ additional samples of $(\boldsymbol{\tilde{X}}_{j})_{j\in[N]}$ from $\mathbb{E}_{\text{risk neutral measure}(\boldsymbol{\theta})}$, directly according to its marginal distribution. Importantly, these samples are independent of the previous data $(\boldsymbol{Y}_{i},\boldsymbol{X}_{i})_{i\in[n]}$.
  4. (4)
@@
    $PEMC:=\frac{1}{n}\sum_{i=1}^{n}(f(\boldsymbol{Y}_{i})-g(\boldsymbol{X}_{i}))+\frac{1}{N}\sum_{j=1}^{N}g(\boldsymbol{\tilde{X}}_{j})$ (6)
For notational simplicity, we have written ${g}(\boldsymbol{X})$ instead of ${g}(\boldsymbol{\theta}_{\text{model}},\boldsymbol{\theta}_{\text{simulation}},\boldsymbol{\theta}_{\text{payoff}},\boldsymbol{X})$, and $f(\boldsymbol{Y})$ instead of $f_{\boldsymbol{\theta}_{\text{payoff}}}(\boldsymbol{Y})$. A key aspect of this procedure is the relationship between $N$ and $n$. Typically, we choose $N$ to be one, or several orders of magnitude larger than $n$. The optimal choice of $N$ versus $n$, will be discussed Section 4. The PEMC evaluation step is summarized in Algorithm 2.
1:procedure Evaluation($\boldsymbol{\theta},n,N,g$)
2:     Generate i.i.d. $(\boldsymbol{Y}_{i},\boldsymbol{X}_{i})_{i\in[n]}$ from $\mathbb{E}_{\text{risk neutral measure}(\boldsymbol{\theta})}$
3:     Generate i.i.d. $(\boldsymbol{\tilde{X}}_{j})_{j\in[N]}$ independently from its marginal in $\mathbb{E}_{\text{risk neutral measure}(\boldsymbol{\theta})}$
4:     Compute $PEMC:=\frac{1}{n}\sum_{i=1}^{n}(f(\boldsymbol{Y}_{i})-g(\boldsymbol{X}_{i}))+\frac{1}{N}\sum_{j=1}^{N}g(\boldsymbol{\tilde{X}}_{j})$
@@
3.5. PEMC: An Example on Asian Options
We now return to our example on arithmetic Asian option under the Heston model:
@@ 3.5. PEMC: An Example on Asian Options
where $\boldsymbol{\theta_{\text{payoff}}}$ specifies the strike level $K$, the sampling frequency $n_{D}$ and the observation dates $\{t_{i}\}_{i\in[n_{D}]}$. The Heston model is described in (1) which requires $\boldsymbol{\theta}_{\text{model}}:=(r,\eta,\delta,\rho,\kappa)\in\mathbb{R}^{5}$. For this example, let’s say we use a neural network (NN) as the prediction model. Then, the PEMC framework could proceed as follows:
  1. (1)
    Define the parameter space $\boldsymbol{\Theta}$ that encompasses realizations of
    $\boldsymbol{\theta}:=(\underbrace{r,\eta,\delta,\rho,\kappa}_{\boldsymbol{\theta}_{\text{model}}},\underbrace{S_{0},\nu_{0},\Delta t,T}_{\boldsymbol{\theta}_{\text{simulation}}},\underbrace{K,n_{D},\{t_{i}\}_{i\in[n_{D}]}}_{\boldsymbol{\theta}_{\text{payoff}}})$
    which are of practical interest.
  2. (2)
    Uniformly sample $\boldsymbol{\theta}$ from $\boldsymbol{\Theta}$. This step is straightforward if the parameter space is a Cartesian product of intervals. For each sampled $\boldsymbol{\theta}$, generate process $\boldsymbol{Y}:=(S_{t},\nu_{t})_{t}$ using a discretization scheme with step size $\Delta t$, with the Heston model specified by $\boldsymbol{\theta}$.
  3. (3)
    During the sampling of $\boldsymbol{Y}$, we collect $\boldsymbol{X}$ to be the sum of Brownian increment
@@
    simulated during each step of the Euler scheme. This makes the marginal of $\boldsymbol{X}$ simply a two-dimensional Gaussian: $\frac{\boldsymbol{X}}{\sqrt{T}}\sim\mathcal{N}\Big{(}\mathbf{0},[\begin{smallmatrix}1&\rho\\ \rho&1\end{smallmatrix}]\Big{)}$.
  4. (4)
@@
  6. (6)
    Train a NN with weights $\boldsymbol{w}$ to minimize the MSE loss:
    $\min_{{\boldsymbol{w}}}\frac{1}{N_{\text{train}}}\sum_{i\in[N_{\text{train}}]}(\text{label}_{i}-NN_{\boldsymbol{w}}(\text{feature}_{i}))^{2}$
@@
    $NN_{\boldsymbol{\hat{w}}}\approx\mathbb{E}\bigg{[}(\frac{1}{n_{D}}\sum_{i=1}^{n_{D}}S_{t_{i}}-K)^{+}\bigg{|}{r,\eta,\delta,\rho,\kappa},{S_{0},\nu_{0},\Delta t,T},{K,n_{D},\{t_{i}\}_{i\in[n_{D}]}},W_{T}^{S},W_{T}^{\nu}\bigg{]}.$
  7. (7)
    At evaluation, a specific $\boldsymbol{\theta}=({r,\eta,\delta,\rho,\kappa},{S_{0},\nu_{0},\Delta t,T},{K,n_{D},\{t_{i}\}_{i\in[n_{D}]}},W_{T}^{S},W_{T}^{\nu})\in\Theta$ is given, we generate $n$ paired samples $(\textit{label}_{i},\textit{feature}_{i})_{i\in[n]}$ as in (4). We also generate $N$ i.i.d. $\sqrt{T}\mathcal{N}\Big{(}\mathbf{0},[\begin{smallmatrix}1&\rho\\ \rho&1\end{smallmatrix}]\Big{)}$ samples and store them as $(\tilde{\textit{feature}})_{j\in[N]}$.
  8. (8)
    Output $PEMC:=\frac{1}{n}\sum_{i=1}^{n}(\textit{label}_{i}-NN_{\boldsymbol{\hat{w}}}(\textit{feature}_{i}))+\frac{1}{N}\sum_{j=1}^{N}NN_{\boldsymbol{\hat{w}}}(\tilde{\textit{feature}}_{j})$
3.6. Implementation Details in PEMC
@@
3.6.1. Parameter Space $\Theta$
@@
When determining the parameter space $\boldsymbol{\Theta}$, a key consideration is how frequently the model will be updated. To ensure the performance of PEMC, it is desirable that the training data encompass all practical scenarios the model is expected to encounter within the designated update period. For example, if one is calibrating a Heston model (1) for certain SPY ETF, on a given trading day, the calibration yields a specific set of parameters: $(S_{0},\nu_{0},r,\eta,\delta,\rho,\kappa)=(520,20\%,4.5\%,0.04,0.3,-0.7,0.2)$ and the user of PEMC chooses to update the prediction model in PEMC once a month, then, based on historical data and market conditions, financial practitioners could make reasonable guesses or confidence intervals about the bounds within which these parameters are likely to fluctuate over the coming month.
@@
This process is conceptually similar to the selection of uncertainty sets in robust optimization [16] or distributionally robust optimization [50], where the goal is to include realizations of parameters with high likelihood. However, unlike robust optimization approaches, which often favor data-dependent uncertainty sets, our focus here is also guided by the expertise of financial engineers. The training parameter space is designed to reflect realistic and practical scenarios derived from domain knowledge, rather than being strictly driven by statistical guarantees. In applications, we construct reasonable $\boldsymbol{\Theta}$ for practical models, such as forward curves in HJM modeling [60] or local stochastic volatility 2D grids [56], among others.
3.6.2. The Design of $\Large{X}$: Feature Engineering
@@
In the example, we designed the feature $\boldsymbol{X}$ as the sum of two correlated Brownian increments. As discussed in Section 3.3, the choice was motivated by two key properties:
@@
  • Predictability: The sum of Brownian motion increments explains away a portion of the variance of the SDE.
@@
  • Marginal simulation: The sum of two Brownian increments follows a correlated Gaussian distribution, which can be generated efficiently.
@@
Given these considerations, it is tempting to employ Gaussian sums as a generic recipe for the majority of option pricing problems handled by PEMC, particularly since these problems commonly involve stochastic differential equations driven by Brownian motion. Indeed, this is precisely the approach we adopt in our applications, demonstrating its robustness and effectiveness. In cases involving long simulation paths, one might consider segmenting the sum into smaller increments, though we observed that a single increment usually suffices to achieve significant variance reduction. However, the choice of Brownian increments, while effective, is certainly not the only one. As with most ML tasks, feature engineering that leverages domain expertise can yield better results. We illustrate with one such example, the pricing of a floating strike lookback call option, under the Heston model(1). The option payoff is given by:
@@
This approximation leads to the expression: $f(\{S_{t}\}_{t})\approx S_{0}e^{(r-\nu/2)T+\sqrt{\nu}[\boldsymbol{X}]_{1}}-S_{0}e^{[\boldsymbol{X}]_{2}}$, which could serve as a potentially effective approximation of $f(\{S_{t}\}_{t})$. This corresponds directly to Property 1 in the construction of $\boldsymbol{X}$ in 3.3. For Property 2, the marginal distribution of $\boldsymbol{X}$ permits efficient and parallelizable simulation. In particular, applying Girsanov’s Theorem enables one to derive a closed-form expression for the joint density of $([\boldsymbol{X}]_{1},[\boldsymbol{X}]_{2})$ [78]. This result makes it straightforward to sample these components directly.
3.6.3. Eliminating Data Storage Overheads
In typical ML workflows, training datasets—often costly, scarce, and carefully curated—are stored and reused extensively. By contrast, within the PEMC framework, the nature of data generation during the prediction model training in Algorithm 1, actually does not require extensive data storage. This difference stems from two main considerations. First, the volume of training data required to achieve a well-performing prediction model can be large. Storing all of it would be both expensive and unnecessary. In PEMC, the training data can be produced directly via Monte Carlo simulation, ensuring an effectively unlimited supply. Second, this flexibility allows for a more efficient workflow. Data can be generated “on the fly” and processed in streaming fashion. For example, to train on a large number of $N_{\text{train}}$ samples, one could iteratively produce small batches, train the model on these batches, and discard them after training. In this manner, we have successfully trained models in the applications using data on the order of $10^{7}\textendash 10^{8}$.
4. Analysis and Performance
In this section, we provide a detailed and comprehensive analysis of the PEMC estimator. All properties of PEMC discussed here pertain to the evaluation stage, where $\boldsymbol{\theta}$ is arbitrary but fixed. Henceforth, we omit the explicit dependence on $\boldsymbol{\theta}$ in the notation (e.g., writing $\boldsymbol{Y}$ in place of $\boldsymbol{Y}(\boldsymbol{\theta})$). We begin by establishing the unbiasedness of our estimator.
4.1. Bias Analysis
Theorem 1 (Unbiasedness).